# A Compact Low-level RF Control System for Advanced Concept Compact Electron Linear Accelerator

### v1

@@ IV.1 Loopback Test of the Compact LLRF Prototype
The average values of magnitude measured with DAC amplitude at 8000 and 32000 are approximately 186 and 752, which follows the desired linear trend. There is not obvious common drifts or oscillations for both magnitude and phase values at the two DAC amplitude levels. The magnitude fluctuation on the pulse top is measured by the percentage of standard deviation respect to average over the flat top. The fluctuation for 8,000 and 32,000 DAC amplitude levels are 0.8% and 0.34%. The phase fluctuation is directly measured with the standard deviation levels, which are 0.47°  and 0.37°  for the DAC levels of 8,000 and 32,000 respectively. Operating the DAC and ADC in higher range benefits the fluctuation levels for both magnitude and phase values within a single pulse. The amplification or attenuation levels in the final LLRF should be optimized with respect to the input and output ranges of ADCs an DACs. Even though the RF signal used in this test is on the lower side of the ADC, the flatness levels achieved are still considerably better than the 1% and 1°  requirement. That indicates the compact LLRF prototype can generate and measure the RF pulses within the desired fluctuation level.
@@ IV.2 RF Frequency Tuning Scheme Tests
As introduced in Section III.1, the wide operation range of ACCEL and the lack of temperature stabilization made an RF frequency tuning scheme necessary. In this section, the technique for fine frequency tuning will be presented and the tuning routine implemented will be tested with both single-cell and dual-cell ACCEL accelerating structures.
@@ IV.2.1 Single-cavity Structure Test
The first test of RF frequency tuning was performed with the simple test setup shown in Figure 7. The RF output of compact LLRF system generates RF pulses at 60 Hz. The RF pulses are injected to the single-cell structure, which is a prototype cavity for ACCEL, via a circulator. The circulator with bandwidth of 4 to 8 GHz, PACL1704000800A, manufactured by DBwave Technologies LLC has been used for this test. The circulator couples the reflected RF signal from the structure and the signal is connected to one of the RF input channels of the compact LLRF system. The cavity reflection signal is sampled directly by the ADC channel and down converted by the following datapath. The IQ samples collected are converted to magnitude and phase for analysis in this section.
@@
For the first step, the test setup was driven by RF pulses with the RF frequency of 5.712 GHz and measured at the same RF frequency. The magnitude and phase values of the cavity reflection signal are shown in Figure 8. The magnitude of the cavity signal has ripples after the initial rise and remains generally flat until the RF pulse is off. The reflection signal remains high over the pulse duration indicating that the RF power injected to the structure is almost fully reflected as a result of driving the system off the resonant frequency of the cavity. When the structure is driven off resonance within certain threshold, the structure will still be energized, but with an extremely low gain. The single-cell structure has been designed to resonant at 5.712 GHz, but machining errors and ambient temperature can introduce offsets to its resonance. The phase level of the reflection signal remains stable until the RF pulse is off and there is a steep linear ramp after the RF is off. The reflection after the RF pulse is off is dominated by the radiation from the structure cavity. The structure radiates the RF energy at its resonance frequency. When the RF frequency and the down conversion frequency are different, it will be reflected as a linear ramp on the phase level. The offset between the structure resonance and the operation RF frequency of the LLRF system can be calculated with Equation 1.
@@
The phase values has been transferred to software and frequency offset has been computed with the slope of phase ramp after the RF is off. The resonant frequency of the single-cell structure computed by using Equation 1 is approximately 2.08 MHz higher than 5.712 GHz. Then the RF frequency of the LLRF has been adjusted to the resonance frequency of the cavity and the procedure above has been repeated. The magnitude and phase levels of the cavity reflection signal after frequency tuning are shown in Figure 9. The magnitude of the reflection ramps down after the first spike and that demonstrates the field filling process in the accelerating structure. The cavity reflection reaches the lowest magnitude when the filling process is completed. As there is no beam and the structure is overcoupled, the second peak on the magnitude of reflection is higher than the first one. The filling and dissipation processes also reflect on the phase level. The phase has been reversed when the structure is fully filled. The magnitude and phase level before the RF pulse is switched off shows that the RF signal can be successfully injected to the structure and established field to fill the structure.
After the RF pulse is off, the magnitude ramps down, which shows the stored RF power radiating and dissipating in the structure. The phase level remains almost flat after the pulse is off, which means the operation frequency of the LLRF is close to the resonance frequency of the structure. That signifies that the RF frequency tuning has been successful for the single-cell structure with a high precision.
@@ IV.2.2 Dual-cavity Structure Test
In the final design of ACCEL, the full accelerating structure will be composed by 26 dual-cell structures. After the RF frequency tuning scheme has been verified to be functional with the single-cell structure, the test has been extended to the dual-cell structure with a higher power SSA. Figure 10 shows the test setup with the dual-cavity structure. The RF pulse generated by the compact LLRF system is amplified by a low noise amplifier, as the output power of the compact LLRF is consider too low to drive the SSA. The SSA amplifies the resultant RF pulse again and the peak power is up to 300 W. The forward RF power is measured at the output of the SSA via a coupler. The output of the SSA is injected to a dual-cell structure via an adapter. The reflection signal is measured at the reflection port via a load and an adapter. Both the forward and reflection signals are attenuated and looped back via the RF input measurement channels.
@@
The test was performed again after the operation RF frequency of the compact LLRF system is tuned to the structure resonant frequency and the magnitude and phase levels of the cavity reflection signal are shown in Figure 12. The magnitude of the reflection signal shows a different field filling process compared with the single-cell structure. The dual-cell structure has been designed to have a coupling factor ($\beta$) around 4.6 and the magnitude is expected to reach the minimum after the first peak in 100 ns. The magnitude of the cavity reflection shown in the figure reaches the minimum around 100 ns as the radiation from the cavity and reflection from the iris almost cancelled with each other. Then the structure cells are filled for about 500 ns until the reflection signal reaches an steady level. As the structure is designed to be overcoupled without the beam, the second peak after the RF is switched off is approximately 60 % higher than the first peak. The phase remains flat after the RF pulse is off, which indicates that the operating frequency of the prototype ACCEL compact LLRF is extremely close to the resonance of the dual-cell structure. This shows that the RF frequency tuning scheme of LLRF system can successfully measure the resonant frequency of the dual-cell accelerating structure and tune operating frequency to be aligned with the cavity resonance frequency.
@@ IV.3 Compact LLRF Prototype Tests with High-power Test Facility
The RF power at the output of the klystron and the input of water load are measured with directional couplers. Figure 14 shows the magnitude and phase levels of the klystron forward and load forward, which are labeled as ’klystron FWD’ and ’Load’ respectively. The magnitude of the first peak of the load forward signal is scaled to match the first peak of klystron forward signal. The magnitude and phase levels of the load follows the klystron forward power as expected and both have a ramp up about 100 ns after the first peak. The flat top with steady magnitude and phase is approximately 600 ns within the 1 $\mu$s RF pulse.
The magnitude fluctuation is evaluated with the percentage of the standard deviation of the magnitude values in the 600 ns flat top with respect to the average magnitude in the period. The magnitude fluctuation levels for the klystron forward and load forward are 1.60 % and 1.51 % respectively. The phase fluctuation level is measured directly with the standard deviation of the phase values on the flat top. The phase fluctuation levels for the klystron forward and load forward are 1.66°  and 1.98°  respectively. The magnitude and phase fluctuation levels at klystron forward and load forward are consistent, which means the RF power generated by the klystron has been successfully delivered to the load in the test bunker and RF signals are measured with high precision by the compact LLRF platform. The fluctuation levels for both magnitude and phase measured on the flat top are above the 1% and 1°  requirement of ACCEL, which verified the necessity of pulse shape control scheme.
Figure 15 shows the phase levels of klystron forward and load forward signals on the 600 ns flat top. The average phase of each of the signals has been subtracted. There is a linear phase ramp start from approximately -4°  and up to 3°  until the end of the flat top on both of the signal. The ACCEL system requires flat top between 1 to 5 $\mu$s, so the phase drift will be even more significant. The phase drift introduced by the klystron is far higher the ACCEL requirement. The drift need to be compensated by the pulse shape control scheme with phase modulation techniques. As the modulation of the compact LLRF is fully implemented in digital domain, the phase and magnitude modulation can be implemented with high flexibility. We have tested a similar RFSoC based LLRF with of several phase modulation schemes at high power regime. The high-power test demonstrated the capability of generating and measuring phase modulation schemes with high precision and the full results will be published in the near future.
@@ V Conclusion
A compact LLRF system prototype based on RFSoC technology has been designed and developed for the ACCEL program. Compared with conventional LLRF systems, the higher integration level and the direct RF sampling of RFSoC offers distinctive SWaP advantages for the compact LLRF system we proposed for ACCEL. Due to the special operation and application requirements of ACCEL, additional control schemes, such as the automated RF frequency tuning and pulse shape control, are required. The initial control schemes proposed, including: the time sequence, the building blocks and the interface between software and firmware have been introduced. The hardware, firmware, and software of the compact LLRF platform has been designed, implemented and tested, with each block at different levels of completion. The development of the complete prototype will continue in the next phase of the project.
The prototype enabled a range of feasibility studies and experimental tests for different purposes. In the initial loopback test, the compact LLRF system demonstrated magnitude and phase fluctuation levels at 0.34% and 0.37°  on the flat top of a 2 $\mu$s RF pulse. The RF frequency tuning scheme test demonstrates the scheme can measure the resonant frequency of single-cell and dual-cell ACCEL accelerating structure prototypes and tune the operation RF frequency of the LLRF system with high precision. The field filling, reflecting and dissipation processes can be visualized on the reflection signal from the structure before and after the RF frequency is tuned.
@@
There are a range of test and development tasks to be completed before the compact LLRF can be operated with all the control schemes required. The firmware building blocks of the pulse-to-pulse feedback control scheme have been implemented, but not yet tested in a full-firmware implementation. Due to the limits on the schedule, the high-power test has been performed at a relative low power level and without an ACCEL accelerating structure. A high-power test with the ACCEL structure and the compact LLRF platform should be performed in the next phase of the project.

### v2

@@ IV.1 Loopback Test of the Compact LLRF Prototype
The average values of magnitude measured with DAC amplitude at 8000 and 32000 are approximately 186 and 752, which follows the desired linear trend. There is not obvious common drifts or oscillations for both magnitude and phase values at the two DAC amplitude levels. The magnitude fluctuation on the pulse top is measured by the percentage of standard deviation with respect to average over the flat top. The fluctuation for the 8,000 and 32,000 DAC amplitude levels are 0.8% and 0.34%. Phase fluctuation is measured directly with the standard deviation levels, which are 0.47°  and 0.37°  for the DAC levels of 8,000 and 32,000 respectively. Operating the DAC and ADC in the higher range benefits the fluctuation levels for both magnitude and phase values within a single pulse. The amplification or attenuation levels in the final LLRF system should be optimized with respect to the input and output ranges of the ADCs and DACs. Although the RF signal used in this test is on the lower side of the ADC, the flatness levels achieved are still considerably better than the requirement of 1% and 1°  which indicates that the compact LLRF prototype can generate and measure the RF pulses within the desired fluctuation level.We have analyzed the fluctuations on the pulse top in frequency domain, but no consistent periodic disturbance was identified.
@@ IV.2 RF Frequency Tuning Scheme Tests
As introduced in Section III.1, the ACCEL is required to operate in a wide operating temperature range, and there is no temperature stabilization for the accelerating structure. The resonant frequency of the structure changes with temperature. Therefore, an RF frequency tuning scheme has been designed to track the resonant frequency of the accelerating structure and align the operating RF frequency with the resonance. In this section, the technique for fine frequency tuning will be presented and the tuning routine implemented will be tested with both single-cell and dual-cell ACCEL accelerating structures.
@@ IV.2.1 Single-cavity Structure Test
The first test of RF frequency tuning was performed with the simple test setup shown in Figure 7. The RF output of the compact LLRF system generates RF pulses at 60 Hz. The RF pulses are injected into the single-cell structure, which is a prototype cavity for ACCEL, via a circulator. The circulator with a bandwidth of 4 to 8 GHz, PACL1704000800A, manufactured by DBwave Technologies LLC has been used for this test. The circulator couples the reflected RF signal from the structure, and the signal is connected to one of the RF input channels of the compact LLRF system. The cavity reflection signal is sampled directly by the ADC channel and down converted by the following datapath. The IQ samples collected are converted to magnitude and phase for analysis in this section.
@@
For the first step, the test setup was driven by RF pulses with the RF frequency of 5.712 GHz and measured at the same RF frequency. The magnitude and phase values of the cavity reflection signal are shown in Figure 8. The magnitude of the cavity signal has ripples after the initial rise and remains generally flat until the RF pulse is off. The reflection signal remains high over the pulse duration, indicating that the RF power injected to the structure is almost fully reflected as a result of driving the system off the resonant frequency of the cavity. When the structure is driven off resonance within certain threshold, the structure will still be energized, but with an extremely low gain. The single-cell structure has been designed to resonant at 5.712 GHz, but machining errors and ambient temperature can introduce offsets to its resonance. The phase level of the reflection signal remains stable until the RF pulse is off and there is a steep linear ramp after the RF is off. The reflection after the RF pulse is off is dominated by the radiation from the structure cavity. The structure radiates the RF energy at its resonance frequency. When the RF frequency and the down conversion frequency are different, it will be reflected as a linear ramp on the phase level. The offset between the structure resonance and the operation RF frequency of the LLRF system can be calculated with Equation 1.
@@
The phase values has been transferred to software and frequency offset has been computed with the slope of phase ramp after the RF is off. The resonant frequency of the single-cell structure computed using Equation 1 is approximately 2.08 MHz higher than 5.712 GHz. Then the RF frequency of the LLRF has been adjusted to the resonance frequency of the cavity, and the procedure above has been repeated. The magnitude and phase levels of the cavity reflection signal after frequency tuning are shown in Figure 9. The magnitude of the reflection ramps down after the first spike and that demonstrates the field filling process in the accelerating structure. The cavity reflection reaches the lowest magnitude when the filling process is complete. As there is no beam and the structure is overcoupled, the second peak on the magnitude of reflection is higher than the first one. The filling and dissipation processes also reflect on the phase level. The phase has been reversed when the structure is fully filled. The magnitude and phase levels before the RF pulse is switched off show that the RF signal can be successfully injected to the structure and established field to fill the structure.
After the RF pulse is off, the magnitude ramps down, which shows the stored RF power radiating and dissipating in the structure. The phase level remains almost flat after the pulse is off, which means that the operation frequency of the LLRF is close to the resonance frequency of the structure. That signifies that the RF frequency tuning has been successful for the single-cell structure with high precision.
@@ IV.2.2 Dual-cavity Structure Test
In the final design of ACCEL, the full accelerating structure will be composed by 26 dual-cell structures. After the RF frequency tuning scheme has been verified to be functional with the single-cell structure, the test has been extended to the dual-cell structure with a higher power SSA. Figure 10 shows the test setup with the dual-cavity structure. The RF pulse generated by the compact LLRF system is amplified by a low noise amplifier, as the output power of the compact LLRF is too low to drive the SSA. The SSA amplifies the resultant RF pulse again, and the peak power is up to 300 W. The forward RF power is measured at the output of the SSA via a coupler. The output of the SSA is injected into a dual-cell structure via an adapter. The reflection signal is measured at the reflection port through a load and an adapter. Both the forward and the reflection signals are attenuated and looped back via the RF input measurement channels.
@@
The test was performed again after the operation RF frequency of the compact LLRF system is tuned to the structure resonant frequency and the magnitude and phase levels of the cavity reflection signal are shown in Figure 12. The magnitude of the reflection signal shows a different field filling process compared with the single-cell structure. The dual-cell structure has been designed to have a coupling factor ($\beta$) around 4.6 and the magnitude is expected to reach the minimum after the first peak in 100 ns. The magnitude of the cavity reflection shown in the figure reaches the minimum value around 100 ns as the radiation from the cavity and reflection from the iris almost canceled with each other. Then the structure cells are filled for about 500 ns until the reflection signal reaches an steady level. As the structure is designed to be overcoupled without the beam, the second peak after the RF is switched off is approximately 60 % higher than the first peak. The phase remains flat after the RF pulse is off, indicating that the operating frequency of the prototype ACCEL compact LLRF is extremely close to the resonance of the dual-cell structure. This shows that the RF frequency tuning scheme of the LLRF system can successfully measure the resonant frequency of the dual-cell accelerating structure and the tune operating frequency to be aligned with the cavity resonance frequency.
@@ IV.3 Compact LLRF Prototype Tests with High-power Test Facility
The RF power at the output of the klystron and the input of water load are measured with directional couplers. Figure 14 shows the magnitude and phase levels of the klystron forward and load forward, which are labeled as ’klystron FWD’ and ’Load’ respectively. The magnitude of the first peak of the load forward signal is scaled to match the first peak of the klystron forward signal. The magnitude and phase levels of the load follow the klystron forward power as expected and both have a ramp up about 100 ns after the first peak. The flat top with steady magnitude and phase is approximately 600 ns within the 1 $\mu$s RF pulse.
The magnitude fluctuation is evaluated with the percentage of the standard deviation of the magnitude values in the 600 ns flat top with respect to the average magnitude in the period. The magnitude fluctuation levels for the klystron forward and load forward are 1.60 % and 1.51 % respectively. The phase fluctuation level is measured directly with the standard deviation of the phase values on the flat top. The phase fluctuation levels for the klystron forward and load forward are 1.66°  and 1.98°  respectively. The magnitude and phase fluctuation levels at klystron forward and load forward are consistent, which means the RF power generated by the klystron has been successfully delivered to the load in the test bunker and RF signals are measured with high precision by the compact LLRF platform. The fluctuation levels for both magnitude and phase measured on the flat top are above the 1% and 1°  requirement of ACCEL, which verified the necessity of the pulse shape control scheme.
Figure 15 shows the phase levels of the klystron forward and load forward signals on the 600 ns flat top. The average phase of each of the signals has been subtracted. There is a linear phase ramp starting from approximately -4°  and up to 3°  until the end of the flat top on both of the signal. The ACCEL system requires a flat top between 1 to 5 $\mu$s, so the phase drift will be even more significant. The phase drift introduced by the klystron is much higher the ACCEL requirement. The drift needs to be compensated for by the pulse shape control scheme with phase modulation techniques. As the modulation of the compact LLRF is fully implemented in the digital domain, the phase and magnitude modulation can be implemented with high flexibility. We have tested a similar RFSoC based LLRF with several phase modulation schemes at high power regime. The high-power test demonstrated the capability to generate and measure phase modulation schemes with high precision, and the full results will be published in the near future.
@@ V Conclusion
A compact LLRF system prototype based on RFSoC technology has been designed and developed for the ACCEL program. Compared with conventional LLRF systems, the higher integration level and direct RF sampling of RFSoC offer distinctive SWaP advantages for the compact LLRF system we proposed for ACCEL. Due to the special operation and application requirements of ACCEL, additional control schemes, such as automated RF frequency tuning and pulse shape control, are required. The initial control schemes proposed, including: the time sequence, the building blocks, and the interface between software and firmware have been introduced. The hardware, firmware and software of the compact LLRF platform have been designed, implemented and tested, with each block at different levels of completion. The development of the complete prototype will continue in the next phase of the project.
The prototype enabled a range of feasibility studies and experimental tests for different purposes. In the initial loopback test, the compact LLRF system demonstrated magnitude and phase fluctuation levels at 0.34% and 0.37°  on the flat top of a 2 $\mu$s RF pulse. The RF frequency tuning scheme test demonstrates that the scheme can measure the resonant frequency of single-cell and dual-cell ACCEL accelerating structure prototypes and tune the operation RF frequency of the LLRF system with high precision. The field filling, reflecting, and dissipation processes can be visualized on the reflection signal from the structure before and after the RF frequency is tuned.
@@
There are a range of test and development tasks that must be completed before the compact LLRF can be operated with all the required control schemes. The firmware building blocks of the pulse-to-pulse feedback control scheme have been implemented, but not yet tested in a full-firmware implementation. In this phase of the project, we did not have enough resources to perform high power tests with the ACCEL prototype and more development effort is required to demonstrate feedback correction. For the next phase of the project, we plan to start the development by demonstrating a simple loopback that drives the structure, measures the RF field, and adjusts the pulse with respect to a user defined set-point. Once we have some test data, the feedback loop design can be tuned based on it. Then the complete compact LLRF platform can be tested with the ACCEL prototype structure in higher power regime.